# Transposed Poisson Structures on Two Nambu $3$-Lie Algebras

### v1

@@ 2 Preliminaries
Let $G$ be an abelian group, $L=\bigoplus\limits_{g\in G}{L_{g}}$ a $G$-graded 33-Lie algebra, we say that a $\frac{1}{3}$-derivation $\varphi$ has degree $g$ (and denoted by $deg(\varphi)$) if $\varphi(L_{h})=L_{g+h}$. Let $V$ be a $G$-graded $L$-module, $\mathrm{Der}_{\frac{1}{3}}(L,V)$ denote the space of $\frac{1}{3}$-derivations and write $\mathrm{Der}_{\frac{1}{3}}(L,V)_{g}=\{\varphi\in\mathrm{Der}_{\frac{1}{3}}(L,V)|deg(\varphi)=g\}$. We give the following theorem, it is crucial in our work.
@@
Let $G$ be an abelian group, $L=\bigoplus\limits_{g\in G}{L_{g}}$ be a finitely generated $G$-graded 33-Lie algebra. Let $V=\bigoplus\limits_{g\in G}{V_{g}}$ be a $G$-graded $L$-module. Then
$\mathrm{Der}_{\frac{1}{3}}(L,V)=\bigoplus_{g\in G}{\mathrm{Der}_{\frac{1}{3}}(L,V)_{g}}.$
@@
For any $g\in G$, we let $p_{g}:L\to L_{g}$ and $\pi_{g}:V\to V_{g}$ denote the canonical projections. Since $L$ is finitely generated, there is a finite subset $S\subset L$ generating $L$. Let $\varphi:L\to V$ be a $\frac{1}{3}$-derivation. Then there are finite sets $Q,R\subset G$ such that
$S\subset\sum_{g\in Q}{L_{g}}\quad and\quad\varphi(S)\subset\sum_{g\in R}{V_{g}}.$ (2.4)
For $g\in G$ put $\varphi_{g}:=\sum\limits_{h\in G}{\pi_{g+h}\mathbin{\circ}\varphi\mathbin{\circ}p_{h}}$. Since for $x_{r}\in L_{r}$, $x_{s}\in L_{s}$ and $x_{t}\in L_{t}$. we have
$\varphi_{g}([x_{r},x_{s},x_{t}])=$ $\sum_{h\in G}{\pi_{g+h}\mathbin{\circ}\varphi\mathbin{\circ}p_{h}}([x_{r},x_{s},x_{t}])$
@@
it follows that $\varphi_{g}$ is contained in $\mathrm{Der}_{\frac{1}{3}}(L,V)_{g}$.
@@
$=$ $\sum_{g\in R}{\sum_{h\in Q}{\pi_{g}\mathbin{\circ}\varphi\mathbin{\circ}p_{h}(y)}}$
$=$ $\sum_{h\in Q}{(\sum_{g\in R}{\pi_{(g-h)+h}\mathbin{\circ}\varphi\mathbin{\circ}p_{h}(y)})}$
$=$ $\sum_{h\in Q}{(\sum_{q\in T}{\pi_{q+h}\mathbin{\circ}\varphi\mathbin{\circ}p_{h}(y)})}$
$=$ $\sum_{q\in T}{\sum_{h\in Q}{\pi_{q+h}\mathbin{\circ}\varphi\mathbin{\circ}p_{h}(y)}}$
$=$ $\sum_{q\in T}{\sum_{h\in G}{\pi_{q+h}\mathbin{\circ}\varphi\mathbin{\circ}p_{h}(y)}}$
@@ 3 Transposed Poisson structures on 3-Lie algebra $A_{\omega}^{\delta}$
The $\frac{1}{3}$-derivation algebra of $A_{\omega}^{\delta}$ is
@@ 4 Transposed Poisson structures on 3-Lie algebra $A_{f,k}$
[4] Let $f:A\to{\mathbb{F}}$ be a nonzero linear function with $f(L_{r})=0$ for all $r\in{\mathbb{Z}}$, then for any integer $k$, $(A_{f,k},[\cdot,\cdot,\cdot])$ is a canonical Nambu 3-Lie algebra, where $\cdot$$\cdot,\cdot$] is defined by

### v2

@@ 2 Preliminaries
Let $G$ be an abelian group, $L=\bigoplus\limits_{g\in G}{L_{g}}$ a $G$-graded 33-Lie algebra, we say that a $\frac{1}{3}$-derivation $\varphi$ has degree $g$ (and denoted by $deg(\varphi)$) if $\varphi(L_{h})=L_{g+h}$. Denote by $\mathrm{Der}_{\frac{1}{3}}(L)$ the vector space of all $\frac{1}{3}$-derivations and write $\mathrm{Der}_{\frac{1}{3}}(L)_{g}=\{\varphi\in\mathrm{Der}_{\frac{1}{3}}(L)|deg(\varphi)=g\}$. We give the following theorem, it is crucial in our work.
@@
Let $G$ be an abelian group, $L=\bigoplus\limits_{g\in G}{L_{g}}$ be a finitely generated $G$-graded 33-Lie algebra. Then
$\mathrm{Der}_{\frac{1}{3}}(L)=\bigoplus_{g\in G}{\mathrm{Der}_{\frac{1}{3}}(L)_{g}}.$
@@
For any $g\in G$, we let $\pi_{g}:L\to L_{g}$ denote the canonical projection. Since $L$ is finitely generated, there is a finite subset $S\subset L$ generating $L$. Let $\varphi:L\to V$ be a $\frac{1}{3}$-derivation. Then there are finite sets $Q,R\subset G$ such that
$S\subset\sum_{g\in Q}{L_{g}}\quad and\quad\varphi(S)\subset\sum_{g\in R}{L_{g}}.$ (2.4)
For $g\in G$, put $\varphi_{g}:=\sum\limits_{h\in G}{\pi_{g+h}\mathbin{\circ}\varphi\mathbin{\circ}\pi_{h}}$. Since for $x_{r}\in L_{r}$, $x_{s}\in L_{s}$ and $x_{t}\in L_{t}$, we have
$\varphi_{g}([x_{r},x_{s},x_{t}])=$ $\sum_{h\in G}{\pi_{g+h}\mathbin{\circ}\varphi\mathbin{\circ}\pi_{h}}([x_{r},x_{s},x_{t}])$
@@
it follows that $\varphi_{g}$ is contained in $\mathrm{Der}_{\frac{1}{3}}(L)_{g}$.
@@
$=$ $\sum_{g\in R}{\sum_{h\in Q}{\pi_{g}\mathbin{\circ}\varphi\mathbin{\circ}\pi_{h}(y)}}$
$=$ $\sum_{h\in Q}{(\sum_{g\in R}{\pi_{(g-h)+h}\mathbin{\circ}\varphi\mathbin{\circ}\pi_{h}(y)})}$
$=$ $\sum_{h\in Q}{(\sum_{q\in T}{\pi_{q+h}\mathbin{\circ}\varphi\mathbin{\circ}\pi_{h}(y)})}$
$=$ $\sum_{q\in T}{\sum_{h\in Q}{\pi_{q+h}\mathbin{\circ}\varphi\mathbin{\circ}\pi_{h}(y)}}$
$=$ $\sum_{q\in T}{\sum_{h\in G}{\pi_{q+h}\mathbin{\circ}\varphi\mathbin{\circ}\pi_{h}(y)}}$
@@ 3 Transposed Poisson structures on 3-Lie algebra $A_{\omega}^{\delta}$
The space of all $\frac{1}{3}$-derivations of $A_{\omega}^{\delta}$ is
@@ 4 Transposed Poisson structures on 3-Lie algebra $A_{f,k}$
[4] Let $f:A\to{\mathbb{F}}$ be a linear function with $f(L_{r})=0$ for all $r\in{\mathbb{Z}}$, and there exists some $t$ such that $f(M_{t})\neq 0$. Then for any integer $k$, $(A_{f,k},[\cdot,\cdot,\cdot])$ is a canonical Nambu 3-Lie algebra, where $\cdot$$\cdot,\cdot$] is defined by